# Neural ordinary differential equations for model order reduction of stiff systems

### v1

@@ Abstract
  • Neural Ordinary Differential Equations (ODEs) represent a significant advancement at the intersection of machine learning and dynamical systems, offering a continuous-time analog to discrete neural networks. Despite their promise, deploying neural ODEs in practical applications often encounters the challenge of stiffness, a condition where rapid variations in some components of the solution demand prohibitively small time steps for explicit solvers. This work addresses the stiffness issue when employing neural ODEs for model order reduction by introducing a suitable reparametrization in time. The considered map is data-driven and it is induced by the adaptive time-stepping of an implicit solver on a reference solution. We show the map produces a nonstiff system that can be cheaply solved with an explicit time integration scheme. The original, stiff, time dynamic is recovered by means of a map learnt by a neural network that connects the state space to the time reparametrization. We validate our method through extensive experiments, demonstrating improvements in efficiency for the neural ODE inference while maintaining robustness and accuracy. The neural network model also showcases good generalization properties for times beyond the training data.
@@ 1 Introduction
In computational science and engineering, the complexity and scale of numerical simulations have grown exponentially, driven by advances in hardware and algorithms. High-fidelity models can now capture intricate details of physical phenomena across diverse domains such as fluid dynamics, structural analysis, climate modeling, and materials science. However, the increasing complexity of these models often comes at a significant computational cost, rendering them impractical for many applications that require repeated model evaluations over a large number of parameter values. Indeed, the computational cost is driven up by the curse of dimensionality [4].
@@
Reduced Order Models (ROMs) have emerged as a powerful solution to this challenge, offering a balance between accuracy and computational efficiency. By compressing the essential features of high-fidelity models into a more manageable form, usually called latent space, ROMs enable the efficient simulation and analysis of complex systems, especially those requiring real-time analysis [53], inverse problems [22, 25, 45, 39], uncertainty quantification [12, 62, 18], shape optimization [46] or optimal control [49, 57, 7].
@@
The field of ROMs has seen substantial growth, providing practitioners with a wide range of techniques to choose from. Each method comes with its own set of advantages and drawbacks, and the optimal choice of ROM often depends on the specific requirements of the application. One of the most well-established paradigms for model order reduction is proper orthogonal decomposition (POD) [33, 6]. A linear projection technique that provides a high compression rate and, for instance, achieves a particularly high level of accuracy in the case of diffusion processes.
@@
If the governing equations of the high-fidelity model are explicitly employed in the ROM, they are called intrusive [56, 32, 5, 2, 11]. Usually, a greedy algorithm or POD is employed to project the high-fidelity model onto a lower-dimensional space. However, for nonlinear models, additional techniques like the (discrete) empirical interpolation method are required to handle the complexities, which can lead to an unfavorable trade-off between accuracy and computational efficiency [3, 15, 17]. Moreover, in [50] it has been shown that for transport-dominated problems, a large latent dimension must be used to achieve reasonable accuracy when using POD, thus limiting their practical use.
@@
In recent years, data-driven methods based on machine learning and deep learning gained large traction due to advances in GPU-based hardware and developments in algorithms and software supporting the artificial intelligence ecosystem. This paradigm is based on the availability of a large quantity of high-fidelity data that can be used to train the model. The approach proves effective if the large amount of resources employed in the offline phase (generation of the solutions and training of the model) is repaid by a large enough number of (cheap) evaluations of the trained model (online phase). These models are called non-intrusive since only knowledge of the data is needed to build them [34, 66]. The field is ever-growing, with a large number of techniques being developed and published. To the best of our knowledge, the most relevant examples of data-driven approaches are the following. Closure models are a hybrid between POD and data-driven approaches where a non-linear term (a neural network) is added to the linear ROM (usually POD-Galerkin) to account for the effect of the unresolved dynamics [67]. In [50, 65, 9, 47, 26], densely connected or convolutional autoencoders have been used to extract the full latent variables. Other techniques have been proposed to enhance the accuracy. For instance, graph neural networks (GNNs) tackle the problem of complicated geometries discretized with unstructured meshes, offering a more natural representation of variables using a graph [54, 52]. Fourier neural operators are a very successful approach that instead exploits the frequency space to enrich the layers of the model [41, 40, 70]. Depending on the considered model, time evolution may be treated just as any parameter of the system or explicitly learned, examples are DeepONets [44, 50], dynamic mode decomposition [60, 8], sparse identification of reduced latent dynamics (SINDy) [16, 10], RNNs [42], LSTMs [65, 21] or Gaussian processes [29].
@@
Another time evolution approach of particular interest is neural ODEs [19, 58, 59, 55]. The core concept of this method is straightforward: it involves learning the right-hand side of a dynamical system with a dense neural network, which can be viewed as the continuous limit of a residual neural network. This approach offers numerous advantages, including a more natural representation of continuous-time data and enhanced parameter and memory efficiency. The primary challenge with neural ODEs is that they are in all effects ODEs that must be solved. This often leads to issues with stiffness, complicating their solution and potentially affecting their stability and accuracy. Stiffness is the condition where there are rapid changes in some components of the solution while other components change slowly. This phenomenon creates numerical challenges when solving the equations, as explicit numerical methods might become inefficient or unstable [69]. Thus, the necessity of employing implicit solvers significantly drives up the cost.
@@
In the present work, we tackle the problem of building a ROM of a stiff system. In this case, the main bottleneck is not represented by the large number of variables of the system but by the stiffness of the system. Indeed, stiffness is usually connected to features of the system, such as the presence of multiple time scales, the spectrum of the Jacobian of the system, or numerical instability, that make building a ROM a challenging task. In this case, employing fixed-time timesteppers is unfeasible since the rapidly varying scales require unreasonably small steps. On the other hand, treating the time like an input of a dense neural network also proves problematic since neural networks are known to perform badly on high frequency [68]. Still, in literature, there can be found several examples of stiff problems solved with machine learning. A possible solution was first proposed in [37], where the authors employ quasi-steady-state assumption (QSSA) to reduce the stiffness of the ODE systems and show that a physic-informed neural network (PINN) then can be successfully applied to the converted non-/mild-stiff systems. This is particularly interesting since the system can be solved with an explicit method. In [38] the authors show that computing a stabilized gradient and suitable scaling of the network outputs enables learning stiff neural ODEs. In [23], PINNs are with Extreme Learning Machine to solve stiff problems. In [1], stiff quantitative systems pharmacology models are accelerated with echo state networks. In [72], an autoencoder is used to produce a reduced order model of a chemical kinetic model for the simulation of a combustion system. Recently, in [28] DeepONets are used to learn the discretized solution of challenging stiff chemical kinetics.
@@
Our strategy is the following, we aim to employ a neural ODE to learn a time-reparametrizated stiff system, where the time map is suitably built so to reduce the stiffness of the problem. The insight is to exploit the advantages of the neural ODEs we previously described while keeping the (online) cost of solving the ODE under control by using an explicit solver. Indeed, our main contribution is to employ the time reparametrization induced by the time-stepping of an implicit solver as a way to reduce the stiffness of the system. An advantage of this procedure is that it is purely data-driven, meaning that it is possible to avoid manually deriving explicit algebraic expressions for the QSS species as it was done in [37]. Once the nonstiff system is solved, the solution is mapped back to the stiff dynamics using a map learnt with a neural network. Particular care is taken in the definition of the map so that it has no dependence on time. Hence, the model maintains good generalization in time since there is only explicit dependence on the state space.
@@
We demonstrate the accuracy and speed of the proposed ROM on a set of five test problems widely used in literature to benchmark stiff ODE solvers [48]. The results are compared with a state-of-the-art Runge-Kutta implicit solver of the Radau II A kind of the fifth order [69].
@@
The remainder of the paper is organized as follows. First, in Section 2 we introduce the mathematical settings and notation for the problems of interest. In particular, in Section 2.4 and 2.5 we review neural networks and the neural ODEs approach, which are the foundations of this work. In Section 3 we introduce our methodology, discussing also the implementation details. In Section 4 we present a comprehensive set of numerical results. Finally, in Section 5 we draw some conclusions, discussing the strengths and limitations of the proposed approach.
@@ 2.1 Parametric systems of ordinary differential equations
$\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{f}(t,\boldsymbol{u};\boldsymbol{\mu}),\qquad t\in(0,T)\\ \boldsymbol{u}(0;\boldsymbol{\mu})=\boldsymbol{u}_{0}(\boldsymbol{\mu}).\end{cases}$ (1)
where $\boldsymbol{\mu}\in\Gamma$ is a vector containing all the parameters belonging to a compact set $\Gamma\subset\mathbb{R}^{N_{\mu}}$, $\boldsymbol{u}:[0,T)\times\Gamma\rightarrow\mathbb{R}^{N_{u}}$ is the parameterized solution, $\boldsymbol{f}:\mathbb{[}0,T)\times\mathbb{R}^{N_{u}}\to\mathbb{R}^{N_{u}}$ is the right-hand side function (encoding the dynamical system), $t\in[0,T)$ is the time variable, $\boldsymbol{u}_{0}:\Gamma\rightarrow\mathbb{R}^{N_{u}}$ is the initial condition, and $\dot{\boldsymbol{u}}$ is the total derivative with respect to time $t$. Without loss of generality, we are considering that the initial time for the system is zero; indeed there always exists a translation in time that brings us to this case.
@@ 2.2 Solving ordinary differential equations
There is a wide literature of methods to solve ODEs [69]. In this work, we concentrate on the well-known Runge-Kutta methods. Namely, let $b_{i},a_{ij},(i,j=1,...,s)$ be real numbers and let $c_{i}$ be defined as
$c_{i}=\sum_{j=1}^{i-1}a_{ij},\quad\forall i=1,...,s.$
Then, we call the $s$-stage Runge-Kutta method the following discretization of Eq. (1) at time $t_{0}+\Delta t$
$\begin{split}&\boldsymbol{k}_{i}=\boldsymbol{f}\left(t_{0}+c_{i}\Delta t,\boldsymbol{u}_{0}+\Delta t\sum_{j=1}^{s}a_{ij}\boldsymbol{k}_{j}\right),\quad i=1,...,s\\ &\boldsymbol{u}_{1}=\boldsymbol{u}_{0}+\Delta t\sum_{i=1}^{s}b_{i}\boldsymbol{k}_{i},\end{split}$ (2)
where $\boldsymbol{u}_{1}=\boldsymbol{u}(t_{0}+\Delta t)$ and $t_{0}=0$. If $a_{ij}=0$ for $i\leq j$, each stage of Eq. (2) depends only on the previous ones, thus the method is explicit. If $a_{ij}=0$ for $i<j$ and at least one $a_{ii}=0$, the method is called diagonally implicit. In all the other cases, it is called implicit.
Of particular interest are embedded Runge-Kutta methods, which automatically adjust the step size to achieve a prescribed tolerance of the local error. The embedded method exploits a second set of scalars $b_{i}^{*},i=1,...,s$ to compute another approximation of the solution of lower order $\boldsymbol{u}^{*}$. Namely, we want to satisfy the component-wise inequality
$|(\boldsymbol{u}_{1})_{i}-(\boldsymbol{u}_{1}^{*})_{i}|\leq\delta_{i},\qquad\delta_{i}=\texttt{atol}+\texttt{rtol}\max(|(\boldsymbol{u}_{1})_{i}|,|(\boldsymbol{u}_{0})_{i}|),\qquad i=1,...,N_{u}$
@@
$\varepsilon=\sqrt{\frac{1}{N_{u}}\sum_{i=1}^{N_{u}}\left(\frac{(\boldsymbol{u}_{1})_{i}-(\boldsymbol{u}_{1}^{*})_{i}}{\delta_{i}}\right)^{2}}$
@@
$\Delta t_{\text{new}}=\Delta t\min(10,\max(0.1,(1/\varepsilon)^{1/(q+1)})),$
@@
Explicit Runge-Kutta methods are generally unsuitable for the solution of stiff equations because their region of absolute stability is small [69]. Indeed, in the present work, we will employ three different Runge-Kutta methods depending on the application. Namely, we use an implicit Runge-Kutta of the Radau II A kind of the fifth order [30] for solving the stiff problems, and we use a Runge-Kutta of the fourth order (in the adaptive and fixed step size variants) for nonstiff problems. Starting from the paper of Butcher [13], it became customary to symbolize method (2) by the a tableau
$\begin{array}[]{c|cccc}c_{1}&a_{11}&a_{12}&\dots&a_{1s}\\ c_{2}&a_{21}&a_{22}&\dots&a_{2s}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ c_{s}&a_{s1}&a_{s2}&\dots&a_{ss}\\ \hline\cr&b_{1}&b_{2}&\dots&b_{s}\\ &b_{1}^{*}&b_{2}^{*}&\dots&b_{s}^{*}\\ \end{array}$
We have the following table for the Radau II A method
@@
Due to the super-convergence of the Radau IIA method (classical order $p=2s-1$) it is not possible to have an embedded method of order $p-1$ without extra cost. By taking a linear combination of $\Delta t\boldsymbol{f}(\boldsymbol{x}_{0},\boldsymbol{u}_{0})$ and the internal stage values $\boldsymbol{k}_{i}$ it is however possible to get an approximation of order $s$, we refer the interested reader to [30] for details. The following tableau is used for explicit Runge-Kutta of the fourth-order

### v2

@@ Abstract
  • Neural Ordinary Differential Equations (ODEs) represent a significant advancement at the intersection of machine learning and dynamical systems, offering a continuous-time analog to discrete neural networks. Despite their promise, deploying neural ODEs in practical applications often encounters the challenge of stiffness, a condition where rapid variations in some components of the solution demand prohibitively small time steps for explicit solvers. This work addresses the stiffness issue when employing neural ODEs for model order reduction by introducing a suitable reparametrization in time. The considered map is data-driven and it is induced by the adaptive time-stepping of an implicit solver on a reference solution. We show that the map produces a nonstiff system that can be cheaply solved with an explicit time integration scheme. The original, stiff, time dynamic is recovered by means of a map learnt by a neural network that connects the state space to the time reparametrization. We validate our method through extensive experiments, demonstrating improvements in efficiency for the neural ODE inference while maintaining robustness and accuracy when compared to an implicit solver applied to the stiff system with the original right-hand side.
@@ 1 Introduction
In computational science and engineering, the complexity and scale of numerical simulations have grown exponentially, driven by advances in hardware and algorithms. High-fidelity models – also called Full Order Models (FOMs) – can now capture intricate details of physical phenomena across diverse domains such as fluid dynamics, structural analysis, climate modeling, and materials science. However, the increasing complexity of these models often comes at a significant computational cost, rendering them impractical for many applications that require repeated model evaluations over a large number of parameter values. Indeed, the computational cost is driven up by the curse of dimensionality [6].
@@
Reduced Order Models (ROMs) have emerged as a powerful solution to this challenge, offering a balance between accuracy and computational efficiency. By compressing the essential features of FOMs into a more manageable form, usually called latent space, ROMs enable the efficient simulation and analysis of complex systems, especially those requiring real-time analysis [59], inverse problems [23, 27, 51, 43], uncertainty quantification [14, 71, 19], shape optimization [52] or optimal control [55, 64, 9].
@@
The field of ROMs has seen substantial growth, providing practitioners with a wide range of techniques to choose from. Each method comes with its own set of advantages and drawbacks, and the optimal choice of a ROM often depends on the specific requirements of the application. Indeed, the term ROMs refers to multiple approaches, which are rather different from one another.
One of the most well-established paradigms for model order reduction is proper orthogonal decomposition (POD) [35, 8], that is, a linear projection technique that provides a high compression rate and, for instance, achieves a particularly high level of accuracy in the case of diffusion processes.
@@
If the governing equations of the high-fidelity model are explicitly employed in the ROM, they are called intrusive [63, 34, 7, 2, 13]. Usually, a greedy algorithm or POD is employed to project the FOM onto a lower-dimensional space. However, for nonlinear models, additional techniques like the (discrete) empirical interpolation method are required to handle the complexities, which can lead to an unfavorable trade-off between accuracy and computational efficiency [5, 16, 18]. Moreover, in [44] it has been shown that for transport-dominated problems, a large latent dimension must be used to achieve reasonable accuracy when using POD, thus limiting their practical use.
@@
In recent years, data-driven methods based on machine learning and deep learning gained large traction due to advances in GPU-based hardware and developments in algorithms and software supporting the artificial intelligence ecosystem. This paradigm is based on the availability of a large quantity of high-fidelity data that can be used to train the model. This approach proves effective if the large amount of resources employed in the offline phase (generation of the solutions and training of the model) is repaid by a large enough number of (cheap) evaluations of the trained model (online phase). These models are called non-intrusive since only knowledge of the data is needed to build them [36, 76]. We also refer to these kinds of ROMs as surrogate models, given their ability to approximate the behavior of the original FOM from the data. The field is ever-growing, with a large number of techniques being developed and published. To the best of our knowledge, the most relevant examples of data-driven approaches are the following. Closure models are a hybrid between POD and data-driven approaches where a non-linear term (a neural network) is added to the linear ROM (usually POD-Galerkin) to account for the effect of the unresolved dynamics [77]. In [56, 75, 11, 53, 28], densely connected or convolutional autoencoders have been used to extract the full latent variables. Other techniques have been proposed to enhance the accuracy. For instance, graph neural networks (GNNs) tackle the problem of complicated geometries discretized with unstructured meshes, offering a more natural representation of variables using a graph [61, 58]. Fourier neural operators are a very successful approach that instead exploits the frequency space to enrich the layers of the model [46, 45, 80]. Depending on the considered model, time evolution may be treated just as any parameter of the system or explicitly learned, examples are DeepONets [50, 56], dynamic mode decomposition [68, 10], sparse identification of reduced latent dynamics (SINDy) [17, 12], RNNs [47], LSTMs [75, 22] or Gaussian processes [31].
@@
Another time evolution approach of particular interest is based on neural ODEs [20, 65, 66, 62]. The core concept of this method is straightforward: it involves learning the right-hand side of a dynamical system with a dense neural network, which can be viewed as the continuous limit of a residual neural network. This approach offers numerous advantages, including a more natural representation of continuous-time data, and enhanced parameter and memory efficiency. The primary challenge with neural ODEs is that they are in all effects ODEs that must be solved. This often leads to issues with stiffness, complicating their solution and potentially affecting their stability and accuracy. Stiffness is the condition where there are rapid changes in some components of the solution while other components change slowly. This phenomenon creates numerical challenges when solving the equations, as explicit numerical methods might become inefficient or unstable [79]. Thus, the necessity of employing implicit solvers significantly drives up the cost.
@@
In the present work, we tackle the problem of building a ROM of a stiff system. In this case, the main bottleneck is not represented by the large number of variables of the system but by the stiffness of the system. Indeed, there are multiple types of stiffness, which are usually connected to features of the system, such as the presence of multiple time scales, the spectrum of the Jacobian of the system, or numerical instability, that make building a ROM a challenging task. In this case, employing fixed-time timesteppers is unfeasible since the rapidly varying scales require unreasonably small steps. On the other hand, treating time like an input of a dense neural network also proves to be problematic since neural networks are known to perform badly on high frequency [78]. Still, in literature, there can be found several examples of stiff problems solved with machine learning. A possible solution was first proposed in [40], where the authors employ the quasi-steady-state assumption (QSSA) to reduce the stiffness of the ODE systems and show that a physics-informed neural network (PINN) then can be successfully applied to the converted non-/mild-stiff systems. This is particularly interesting since the system can be solved with an explicit method. In [42], the authors show that computing a stabilized gradient and suitable scaling of the network outputs enables learning stiff neural ODEs. In [24], PINNs are used with extreme learning machine to solve stiff problems. In [1], stiff quantitative systems pharmacology models are accelerated with echo state networks. In [82], an autoencoder is used to produce a reduced order model of a chemical kinetic model for the simulation of a combustion system. Recently, in [30] DeepONets are used to learn the discretized solution of challenging stiff chemical kinetics.
@@
Our strategy is to employ a neural ODE-based non-stiff surrogate for stiff dynamics. Namely, we aim to employ a neural ODE to learn a time-reparametrized stiff system, where the time map is suitably built to reduce the stiffness of the problem. The insight is to exploit the advantages of the neural ODEs we previously described while keeping the (online) cost of solving the ODE under control by using an explicit solver. Indeed, our main contribution is to employ the time reparametrization induced by the time-stepping of an implicit solver as a way to reduce the stiffness of the system. An advantage of this procedure is that it is purely data-driven, meaning that it is possible to avoid manually deriving explicit algebraic expressions for the QSS species as it was done in [40]. On the other hand, a drawback of our approach is that it is only worthwhile for ROMs: the expensive phase of gathering data for building the time mapping must be offset by a cheap online phase. Our technique can also be interpreted as a way to embed adaptive (implicit) time-stepping directly into the neural ODE. Thus, we can avoid backpropagation through the implicit solver, which incurs a cubic cost with respect to the number of neural network parameters [42], thereby greatly reducing the computational cost of training the neural ODE. Our reparametrization strategy is also connected with [72, 4, 41], where the authors leverage spatial mappings for adaptive mesh refining to improve computational efficiency and to enhance the representation of nonlinear dynamics in complex systems. Moreover, recent work [49] has demonstrated that a structure-preserving NODE with linear/nonlinear splitting can achieve long-term stability on stiff problems comparable to implicit methods. Once the nonstiff system is solved, the solution is mapped back to the stiff dynamics using a map learnt with a neural network. The neural network is trained once with the neural ODE and shows good generalization properties to parameters outside the training dataset. Particular care is taken in the definition of the map so that it has no dependence on time. Hence, the model maintains good generalization in time for periodic systems since there is only explicit dependence on the state space.
Our technique is designed especially for problems where stiffness can be traced back to the case of large eigenvalues with a negative real part. Indeed, another largely studied type of stiffness is high-frequency stiffness, which corresponds to the case of reasonably large eigenvalues but very high frequency [60]. In these cases, the system features a fast and a slow solution, where the fast solution must be solved accurately to track the slow one. While reparametrization approaches have been proposed [39], they have significant limitations since they need to resolve the fast frequency [3, 67]. Thus, this kind of stiffness remains a challenge for our approach and an area of ongoing research. Indeed, other authors have explored structure-preserving machine learning techniques, such as fast-slow neural networks, to bypass the direct resolution of fast transients in singularly perturbed systems [69].
@@
We demonstrate the accuracy and speed of the proposed ROM on a set of five test problems, where explicit solvers usually fail, widely used in literature to benchmark stiff ODE solvers [54]. The results are compared with a state-of-the-art Runge-Kutta implicit solver of the Radau II A kind of the fifth order [79] applied to the stiff system with the original right-hand side.
@@
The remainder of the paper is organized as follows. First, in Section 2 we introduce the mathematical settings and notation for the problems of interest. In particular, in Section 2.4 we review the neural ODEs approach, which is the foundation of this work. In Section 3 we introduce our methodology, discussing also the implementation details. In Section 4 we present a comprehensive set of numerical results. Finally, in Section 5 we draw some conclusions, discussing the strengths and limitations of the proposed approach.
@@ 2.1 Parametric systems of ordinary differential equations
$\begin{cases}\dot{\boldsymbol{u}}(t;\boldsymbol{\mu})=\boldsymbol{f}(t,\boldsymbol{u};\boldsymbol{\mu}),\qquad t\in(0,T]\\ \boldsymbol{u}(0;\boldsymbol{\mu})=\boldsymbol{u}_{0}(\boldsymbol{\mu}).\end{cases}$ (1)
where $\boldsymbol{\mu}$ is a vector containing all the parameters of the system, which belongs to a compact set $\Gamma\subset\mathbb{R}^{N_{\mu}}$, $\boldsymbol{u}:[0,T]\times\Gamma\rightarrow\mathbb{R}^{N_{u}}$ is the parameterized solution, $\boldsymbol{f}:\mathbb{[}0,T]\times\mathbb{R}^{N_{u}}\times\Gamma\to\mathbb{R}^{N_{u}}$ is the right-hand side function (encoding the dynamical system), $t\in[0,T]$ is the time variable, $\boldsymbol{u}_{0}:\Gamma\rightarrow\mathbb{R}^{N_{u}}$ is the initial condition, and $\dot{\boldsymbol{u}}$ is the total derivative with respect to time $t$. Without loss of generality, we are considering that the initial time for the system is zero; indeed there always exists a translation in time that brings us to this case.
@@ 2.2 Solving ordinary differential equations
There is a wide literature of methods to solve ODEs [79]. In this work, we concentrate on the well-known Runge-Kutta methods. Let $r\in\mathbb{N}$ s.t. $r\geq 1$, let $b_{i},a_{ij},(i,j=1,...,r)$ be real numbers and let $c_{i}$ be defined as
$c_{i}=\sum_{j=1}^{i-1}a_{ij},\quad\forall\,i=1,...,r.$
Then, fixed a timestep $\Delta t>0$ and setting $t^{(n)}=n\Delta t$, we call the $r$-stage Runge-Kutta method the following discretization of Eq. (1)
$\begin{split}&\boldsymbol{k}_{i}=\boldsymbol{f}\left(t^{(n)}+c_{i}\Delta t,\boldsymbol{u}^{(n)}+\Delta t\sum_{j=1}^{r}a_{ij}\boldsymbol{k}_{j}\right),\quad i=1,...,r\\ &\boldsymbol{u}^{(n+1)}=\boldsymbol{u}^{(n)}+\Delta t\sum_{i=1}^{r}b_{i}\boldsymbol{k}_{i},\end{split}$ (2)
$\forall\,n\in\mathbb{N}$ s.t. $t^{(n)}\leq T$. Here, $\boldsymbol{u}^{(n+1)}\approx\boldsymbol{u}(t^{(n)}+\Delta t)$. If $a_{ij}=0$ for $i\leq j$, each stage of Eq. (2) depends only on the previous ones, thus the method is explicit. If $a_{ij}=0$ for $i<j$ and at least one $a_{ii}=0$, the method is called diagonally implicit. In all the other cases, it is called implicit.
Of particular interest are embedded Runge-Kutta methods, which automatically adjust the step size $\Delta t$ to achieve a prescribed tolerance of the local error. The embedded method exploits a second set of scalars $b_{i}^{*},i=1,...,r$ to compute another approximation of the solution of lower order $\boldsymbol{u}^{*}$. Namely, we want to satisfy the component-wise inequality
$|(\boldsymbol{u}^{(n+1)})_{i}-(\boldsymbol{u}^{*(n+1)})_{i}|\leq\delta_{i},\qquad\delta_{i}=\texttt{atol}+\texttt{rtol}\max(|(\boldsymbol{u}^{(n+1)})_{i}|,|(\boldsymbol{u}^{(n)})_{i}|),\qquad i=1,...,N_{u}$
@@
$\varepsilon=\sqrt{\frac{1}{N_{u}}\sum_{i=1}^{N_{u}}\left(\frac{(\boldsymbol{u}^{(n+1)})_{i}-(\boldsymbol{u}^{*(n+1)})_{i}}{\delta_{i}}\right)^{2}}$
@@
$\Delta t^{(n+1)}=\Delta t^{(n)}\min(10,\max(0.1,\varepsilon^{-1/(q+1)})),$
@@
Explicit Runge-Kutta methods are generally unsuitable for the solution of stiff equations because their region of absolute stability is small [79]. Indeed, in the present work, we will employ three different Runge-Kutta methods depending on the application. Namely, we use an implicit Runge-Kutta of the Radau II A kind of the fifth order [32] for solving the stiff problems, and we use a Runge-Kutta of the fourth order (in the adaptive and fixed step size variants) for nonstiff problems. We employ the following table for the Radau II A method
@@
Due to the super-convergence of the Radau II A method (classical order $p=2r-1$) it is not possible to have an embedded method of order $p-1$ without extra cost. By taking a linear combination of $\Delta t\boldsymbol{f}(\boldsymbol{x}_{0},\boldsymbol{u}_{0})$ and the internal stage values $\boldsymbol{k}_{i}$ it is however possible to get an approximation of order $r$, we refer the interested reader to [32] for details. The following tableau is used for explicit Runge-Kutta of the fourth-order [25], also called the Dormand-Prince method or RK45